### v1

# A bound-preserving and conservative enriched Galerkin method for elliptic problems

Abstract
We propose a locally conservative enriched Galerkin scheme that respects the discrete maximum principle of an elliptic problem. To this end, we use a substantial over-penalization of the discrete solution’s jumps to obtain optimal convergence. To avoid the ill-conditioning issues that arise in over-penalized schemes, we introduce an involved splitting approach that separates the system of equations for the discontinuous solution part from the system of equations for the continuous solution part, yielding well-behaved subproblems. We prove the existence of discrete solutions and optimal error estimates, which are validated numerically.
1 Introduction
Physically-consistent numerical methods have become a key focus in the numerical analysis of partial differential equations (PDEs). These methods aim to preserve specific properties of the exact solutions, such as local conservation laws, entropy conditions, discrete maximum principles, or positivity constraints. Such properties are critical for ensuring that numerical simulations of complex phenomena remain physically accurate and stable.
In particular, methods that satisfy the Discrete Maximum Principle (DMP) and monotonicity conditions have been extensively studied within the finite element framework [Cia70, CR73, Kik77, MH85, BE05, BKK08, BBK17] (see also [BJK25] for a recent review on the topic). These properties, and the more general bound preservation criterion, guarantee that numerical solutions adhere to the inherent bounds of the exact solution, often preserving positivity and/or other relevant physical bounds. Ensuring that these numerical methods respect such constraints is crucial, especially for nonlinear PDE models, where violation of bounds can lead to instability or loss of physical accuracy. Examples include reaction-diffusion systems that model concentrations or phase-field models where maximum and minimum bounds constrain solutions.
However, while DMP and monotonicity are substantial structural requirements, methods that focus on bound preservation are less restrictive and, thus, potentially more efficient for various applications. The importance of bound-preserving methods is underscored in scenarios where, e.g., positivity is required to maintain the validity of a physical model. For instance, negative solutions would be non-physical in reaction-diffusion systems or turbulence models, and possibly a stable numerical simulation can be built only by ensuring this requirement, without the need to fulfill the stronger restrictions demanded by the DMP.
In the recent work [BGPV24], the idea of bound-preservation at the nodes was presented for an elliptic reaction-diffusion problem. The core idea involves choosing a baseline discretization of the PDE at hand and combining it with a projection operator that maps finite element functions to their positive (well-behaved) counterparts, ensuring that numerical solutions remain within the required bounds. The method is nonlinear, as anticipated by Godunov’s barrier theorem. For symmetric linear problems, the solution is sought as an orthogonal projection of the exact solution onto the convex set of finite element functions with positive nodal values. This work has been recently extended to convection-dominated problems in [ABP24], using continuous finite element spaces, and to time-dependent problems using discontinuous Galerkin in [BPT25]. A common feature of all the above works is the fact that they can be proven to be equivalent to solving a variational inequality over the constrained convex set (in fact, the method presented in [BPT25] is written directly in such a way, without the link to a stabilized method). This fact prevents the method from being conservative (either locally or globally), a very desirable property in, for example, time-dependent transport problems.
Based on the discussion in the above paragraph, in this work we present a method that preserves the bounds of the continuous solution while at the same time is locally (and globally) conservative. The baseline discretization is the Enriched Galerkin (EG) method, in which the constraints are encoded via Lipschitz-continuous projections. The conservation property is achieved by using enriched Galerkin methods and a stabilization term that disregards element-wise constants. This approach enables the use of simple solvers, such as Richardson-like iterations, thereby avoiding more complex constrained optimization techniques commonly employed in the existing literature [EHS09, BRD+20, MN16]. Moreover, the method achieves near-best approximation properties in suitable error norms.
The enriched Galerkin (EG) methods use the discontinuous Galerkin (DG) bilinear form and a combination of overall continuous and broken polynomial spaces to inherit the polynomial best approximation property from continuous finite elements and other beneficial properties, such as enhanced stability and local mass conservation from discontinuous Galerkin methods, while maintaining fewer degrees of freedom than DG. The first enriched Galerkin method was proposed by Becker et al. [BBHL03], but these methods only gained popularity after reinvention by Lee et al. [LLW16]. Since then, the EG method has been continuously generalized and applied to many problems. Its main strengths stem from its flexibility concerning the polynomial spaces that allow, for example, higher-order or spatially adaptive enrichments [RHA21, RL20]. This flexibility allows us to derive, for instance, locking-free and physics-preserving EG methods for poroelasticity [LY23, YL24] or efficient EG-based solvers for the Stokes problem [YHLA22]. However, this flexibility obstructs bound-preserving approximations even in the lowest-order cases. This issue has been addressed in [KHR20, KH23, KLY24] for hyperbolic problems and will be addressed for elliptic problems in this work.
One feature of our work is the fact that, due to the particular choice of how the bounds are imposed in the discrete solution, the jump terms in the EG formulation need to be over-penalized. This is not a new concept; in fact, several works have explored the over-penalization of DG methods, such as the WOPSIP method [BOS08]. Now, it is a well-known fact that over-penalized methods lead to very ill-conditioned linear systems of equations. Consequently, in [BOS08], a preconditioner has also been introduced to address this issue. To completely circumvent the ill-conditioning, this work exploits the fixed-point strategy used to prove the existence of solutions to propose a decoupled scheme that completely avoids solving ill-conditioned linear systems, thereby eliminating the need to propose appropriate preconditioners. This insight seems to be new even for linear EG methods.
Several numerical experiments demonstrate the effectiveness of the proposed method compared to standard finite element approaches. These experiments show that the method preserves critical structural properties of the exact solution while maintaining computational efficiency. Similar methods that focus on bound preservation, such as truncation-based approaches [LHV13, Kre14, YYZ22], have been previously explored. However, the method presented here offers a fresh perspective on the formulation and numerical solution of bound-preserving problems.
The remainder of this contribution is structured as follows: Section 2 delves into the model problem and the (not bound-preserving) standard enriched Galerkin discretization. In contrast, Section 3 turns this discretization into a bound-preserving one. Section 4 is devoted to proving the existence of discrete and positive solutions, while Section 5 discusses their convergence properties to the analytic solution. Finally, Section 6 demonstrates how our method can be implemented and analyzes the condition of the linear system of equations that needs to be solved internally before a short conclusion wraps up this work.
2 Model problem and baseline discretization
We will adopt standard notation on Sobolev and Lebesgue spaces, aligned with, e.g., [EG21a]. For $D\subseteq\mathbb{R}^{d}$, $d\leq 3$, we denote by $\|\cdot\|_{0,p,D}$ the $L^{p}(D)$-norm; when $p=2$ the subscript $p$ will be omitted and we only write $\|\cdot\|_{0,D}$. In addition, for $s\geq 0$, $p\in[1,\infty]$, we denote by $\|\cdot\|_{s,p,D}$ ($|\cdot|_{s,p,D}$) the norm (seminorm) in $W^{s,p}(D)$; when $p=2$, we define $H^{s}(D)=W^{s,2}(D)$, and again omit the subscript $p$ and only write $\|\cdot\|_{s,D}$ ($|\cdot|_{s,D}$). The following space will also be used repeatedly within the text
$H^{1}_{0}(D)=\left\{v\in H^{1}(D):v=0\;{\rm on}\;\partial D\right\}.$ (1)
Let $\Omega\subset\mathbb{R}^{d},d=2,3$ be a polyhedral, bounded, Lipschitz domain, $f\in L^{2}(\Omega)$, $u_{\mathrm{D}}\in C^{0}(\partial\Omega)$, and let $\epsilon,\mu\in\mathbb{R}^{+}$. We consider the following elliptic reaction-diffusion equation: Find $u\in H^{1}(\Omega)$ such that $u=u_{\mathrm{D}}$ on $\partial\Omega$, and
$\int_{\Omega}\epsilon\nabla u\cdot\nabla v\;\textup{d}\boldsymbol{x}+\int_{\Omega}\mu uv\;\textup{d}\boldsymbol{x}=\int_{\Omega}fv\;\textup{d}\boldsymbol{x},$ (2)
holds for all $v\in H^{1}_{0}(\Omega)$. Using Lax-Milgram’s Lemma (see, e.g., (see, e.g., [EG21b, Lemma 25.2]), this variational problem can be shown to have a unique solution. In addition, using the comparison principle (see [RR06, Corollary 4.4]), then $u$ satisfies
$\|u\|_{0,\infty,\Omega}\leq\tilde{U}:=\max\left\{\frac{\|f\|_{0,\infty,\Omega}}{\mu},\|u_{D}\|_{0,\infty,\partial\Omega}\right\}\,.$ (3)
In addition, if $f\geq 0$ and $u_{D}\geq 0$ we furthermore have that $u\geq 0$. That is, equation (2) respects (physically meaningful) bounds for its solution.
We now describe the baseline discretization of (2). The choice made in this work is the EG method, proposed in [BBHL03, LLW16]. For this, we consider a shape-regular family of conforming simplicial triangulations $({\mathcal{T}_{h}})_{h>0}$. Only to avoid technical diversions and simplify the notations and some of the proofs, we will assume that the family of triangulations is quasi-uniform. For a triangulation ${\mathcal{T}_{h}}$, the set of its facets is denoted by ${\mathcal{F}_{h}}$. For any $T\in{\mathcal{T}_{h}}$, $h_{T}=\operatorname{diam}T$, while $h_{F}=\operatorname{diam}F$ for all $F\in{\mathcal{F}_{h}}$, $h=\max_{T\in{\mathcal{T}_{h}}}h_{T}$, and $h_{\min}=\min_{T\in{\mathcal{T}_{h}}}h_{T}$. In addition, we denote by $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n}$ the internal nodes of ${\mathcal{T}_{h}}$ and for $i\in\{1,\ldots,n\}$ and $T\in{\mathcal{T}_{h}}$ we define the neighborhoods
$\omega_{i}=\{T\in{\mathcal{T}_{h}}:\boldsymbol{x}_{i}\in T\}\quad\textrm{and}\quad\omega_{T}=\{T^{\prime}\in{\mathcal{T}_{h}}:T\cap T^{\prime}\not=\emptyset\}\,.$
The finite element space in the EG method is given by $V_{h}=V_{h}^{1}\oplus V_{h}^{0}$, where
$\tilde{V}_{h}^{1}$ =$=\{v\in C(\overline{\Omega})\colon v|_{T}\in\mathbb{P}_{1}(T)\;\forall T\in{\mathcal{T}_{h}}\},$
$V_{h}^{1}$ $=\tilde{V}_{h}^{1}\cap H^{1}_{0}(\Omega)\,,$
$V_{h}^{0}$ =$=\{v\in L^{2}(\Omega)\colon v|_{T}\in\mathbb{P}_{0}(T)\;\forall T\in{\mathcal{T}_{h}}\}.$
Since $V_{h}\not\subset H^{1}(\Omega)$, the EG method is not conforming. Still, $V_{h}$ is a subset of the broken (linear) polynomial space used in DG methods. Thus, we can use a variant of the DG bilinear and linear forms to define the EG approximate $u_{h}\in V_{h}$. To this end, let us define the jump and average operators. If $F\in{\mathcal{F}_{h}}$ connects two elements $T^{+},T^{-}\in{\mathcal{T}_{h}}$, i.e., $F=T^{+}\cap T^{-}$,the average and jump of $v_{h}\in V_{h}$ are defined by
$\{\!\!\{v_{h}\}\!\!\}=\tfrac{1}{2}v_{h}|_{T^{+}}+\tfrac{1}{2}v_{h}|_{T^{-}}\qquad\text{ and }\qquad[\![v_{h}]\!]=v_{h}|_{T^{+}}\boldsymbol{n}^{+}+v_{h}|_{T^{-}}\boldsymbol{n}^{-},$
where $\boldsymbol{n}^{\pm}$ denotes the outward pointing unit normal of element $T^{\pm}$. Note that the jump turns a scalar function into a vector-valued quantity. If $F\subset\partial\Omega\cap T$ for some $T\in{\mathcal{T}_{h}}$, we use
$\{\!\!\{v_{h}\}\!\!\}=v_{h}\qquad\text{ and }\qquad[\![v_{h}]\!]=v_{h}\boldsymbol{n},$
where $\boldsymbol{n}$ is the outward pointing unit normal to $\partial\Omega$. Moreover, we need $L^{2}$-type scalar products in the mesh’s bulk and skeleton, i.e., concerning ${\mathcal{T}_{h}}$ and ${\mathcal{F}_{h}}$. We define
$(w_{h},v_{h})=(w_{h},v_{h})_{\mathcal{T}_{h}}=\sum_{T\in{\mathcal{T}_{h}}}\int_{T}w_{h}v_{h}\;\textup{d}\boldsymbol{x}\quad\text{ and }\quad\langle\eta_{h},\sigma_{h}\rangle=\langle\eta_{h},\sigma_{h}\rangle_{\mathcal{F}_{h}}=\sum_{F\in{\mathcal{F}_{h}}}\int_{F}\eta_{h}\cdot\sigma_{h}\;\textup{d}\sigma,$
for $w_{h},v_{h}\in V_{h}$ and face-wise defined vectors $\eta_{h},\sigma_{h}$. Importantly, $(\cdot,\cdot)_{\mathcal{T}_{h}}$ trivially extends to expressions like $(\nabla w_{h},\nabla v_{h})_{\mathcal{T}_{h}}$ if we reinterpret the product in its definition. In addition, we define the norms
$\|v_{h}\|^{2}_{{\mathcal{T}_{h}}}=(v_{h},v_{h})_{\mathcal{T}_{h}}\qquad\text{ and }\qquad\|\eta_{h}\|^{2}_{{\mathcal{F}_{h}}}=\sum_{F\in{\mathcal{F}_{h}}}\frac{\epsilon+\mu h_{F}^{2}}{h_{F}}\int_{F}\eta_{h}\cdot\eta_{h}\;\textup{d}\sigma.$ (4)
Using the above notation, the standard EG method is defined as: Find $u_{h}\in V_{h}$ such that
$a_{h}(u_{h},v_{h})=b_{h}(v_{h})\qquad\text{ for all }v_{h}\in V_{h},$ (5a)
where $a_{h}$ denotes the symmetric interior penalty DG bilinear form and $b_{h}$ incorporates the right-hand side $f$ and the boundary data, i.e.,
$a_{h}(w_{h},v_{h})$ $=(\epsilon\nabla w_{h},\nabla v_{h})+(\mu w_{h},v_{h})-\langle\{\!\!\{\epsilon\nabla w_{h}\}\!\!\},[\![v_{h}]\!]\rangle-\langle\{\!\!\{\epsilon\nabla v_{h}\}\!\!\},[\![w_{h}]\!]\rangle$
$\qquad+\left\langle\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\right\rangle,$ (5b)
$b(v_{h})$ $=(f,v_{h})-a_{h}(I_{h}(\tilde{u}_{\textup{D}}),v_{h}).$ (5c)
Above, $\beta\in\mathbb{N}$, $\gamma=\gamma_{0}L_{\Omega}^{\beta-1}$, where $\gamma_{0}>0$ is a non-dimensional stabilization parameter, and $L_{\Omega}$ is a characteristic length of $\Omega$ (for example, it can be taken as the diameter of $\Omega$). In addition, $\tilde{u}_{\textup{D}}$ is an extension of $u_{\textup{D}}$ to the domain $\Omega$ and $\mathcal{I}_{h}:C(\overline{\Omega})\to\tilde{V}_{h}^{1}$ denotes the Lagrange interpolation operator defined in, e.g., [EG21a, Chapter 11].
Remark 1 (Properties of the EG method).
  1. 1.
    In the standard EG (and DG) literature, the choice is $\beta=1$. Nevertheless, it will become clear in the analysis presented below that we will need larger values for $\beta$. So, for the moment, we only state that $\beta\geq 1$.
  2. 2.
    Observing that $V_{h}$ is a subspace of the standard DG space, and that $a_{h}(\cdot,\cdot)$ is an elliptic bilinear form in the DG space, and so it is elliptic in $V_{h}$ as well. More precisely, using standard arguments (see, e.g., [DPE12]), it can be shown that if $\gamma$ is large enough, there exists $c_{0}>0$, independent of $h$ and any physical constant, such that
    $a_{h}(v_{h},v_{h})\geq c_{0}\Big{(}\|\epsilon^{1/2}\nabla v_{h}\|^{2}_{0,\Omega}+\|\mu^{1/2}v_{h}\|_{0,\Omega}^{2}+\sum_{F\in{\mathcal{F}_{h}}}\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}\|[\![v_{h}]\!]\|_{0,F}^{2}\Big{)}=:c_{0}\|v_{h}\|_{a_{h}}^{2}\,.$ (6)
    Then, the well-posedness of (5) follows.
  3. 3.
    The EG method is locally mass conservative, as demonstrated by testing with $v_{h}=1$ in one element $T\in{\mathcal{T}_{h}}$ and $v_{h}=0$ in all other elements $T\in{\mathcal{T}_{h}}$.
  4. 4.
    As it was mentioned earlier, we allow the possibility of taking $\beta\geq 1$. The error analysis requires taking $\beta=4$, which is the value considered in the numerical experiments (although for completeness, results using smaller values will also be presented). If we were to implement the EG method with this negative power, it would significantly affect the condition number of the associated matrix. Motivated by this fact, within the context of our bound-preserving method presented below, we have also introduced an iterative scheme that completely bypasses this issue.
  5. 5.
    Strictly speaking, a standard symmetric interior penalty DG bilinear form would motivate to definition of the stabilization (with a varying $\beta$ scaling) by $\langle\gamma\frac{\epsilon}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\rangle.$ Thus, compared to our definition, it would not include the terms with $\mu$. While this choice will aid in the analysis, it also has a positive impact on stability with respect to the choice of $\gamma$, as discussed in the numerical section.
Remark 2.
In the above method, $\tilde{u}_{\textup{D}}$ is an extension of $u_{\mathrm{D}}$ to the interior of $\Omega$, and $I_{h}$ is the nodal Lagrange interpolation operator. Notably, we have that $u_{h}+I_{h}(\tilde{u}_{\textup{D}})\approx u$. For a standard EG discretization, the extension of $u_{\mathrm{D}}$ is irrelevant, but for the method presented below, it is simpler to consider simply an extension by zero inside the domain, and that will be the choice used herein (see [ABP24, Section 4.1] for details).
3 Positivity-preserving finite elements
As discussed, we require our EG approximate $u_{h}$ to remain within some invariant domain $\mathcal{G}=[a,b]$, but the EG method generally does not guarantee that $u_{h}(\boldsymbol{x})\in\mathcal{G}$ for all $\boldsymbol{x}\in\Omega$. Thus, we need to correct it such that our corrected approximate $u_{h}^{+}\in V_{h}$
  1. 1.
    takes values in $\mathcal{G}$ only,
  2. 2.
    is locally mass conservative and optimally convergent.
To construct $u^{+}_{h}$ let us define a projection operator that attempts to correct a function $v_{h}\in V_{h}$ such that it only takes values in $\mathcal{G}$. The construction of this operator will be based on the split of $V_{h}$ as piecewise linear and piecewise constant parts. So, every $v_{h}\in V_{h}$ is split as
$v_{h}=v^{1}_{h}+v^{0}_{h}\qquad\text{ where }v_{h}\in V_{h},\;v^{1}_{h}\in V^{1}_{h},\;v^{0}_{h}\in V^{0}_{h}\,.$ (7)
Next, we denote by $\phi_{1},\ldots,\phi_{N}$ are the standard basis functions (the ”hat” functions) spanning the space $V_{h}^{1}$. Then, given $w_{h}=w_{h}^{1}+w_{h}^{0}\in V_{h}$ we define the truncation
$P^{w_{h}}_{i}(v_{h})=\max\left[a-\underline{w}_{hi},\;\min\left(v^{1}_{h}(\boldsymbol{x}_{i}),b-\overline{w}_{hi}\right)\right],\qquad\text{ where}$ (8)
$\underline{w}_{hi}=\min\{w^{0}_{h}(\boldsymbol{x})\colon\boldsymbol{x}\in\omega_{i}\}\qquad\text{ and }\qquad\overline{w}_{hi}=\max\{w^{0}_{h}(\boldsymbol{x})\colon\boldsymbol{x}\in\omega_{i}\}.$
Using this truncation, we define the (nonlinear) mapping
$P^{w_{h}}_{h}\colon V_{h}\ni v_{h}\mapsto P^{w_{h}}_{h}(v_{h})=\sum_{i=1}^{N}P^{w_{h}}_{i}(v_{h})\phi_{i}+w^{0}_{h}\in V_{h}.$
Additionally, we introduce the operator $Q^{w_{h}}_{h}(v_{h})=v^{1}_{h}-[P^{w_{h}}_{h}(v_{h})]^{1}$, and emphasize that the superscripts 00 and $1$ always refer to the decomposition according to (7). Finally, we abbreviate
$v^{+}_{h}=v^{1+}_{h}+v^{0}_{h}=P^{v_{h}}_{h}(v_{h})\in V_{h}\qquad\text{ and }\qquad v^{-}_{h}=v^{1-}_{h}=Q^{v_{h}}_{h}(v_{h})\in V^{1}_{h}\subset V_{h},$
where $v^{1+}_{h}=\sum_{i=1}^{N}P^{v_{h}}_{i}(v_{h})\phi_{i}$.
The mapping $P_{h}^{v_{h}}$ is built with the aim of guaranteeing that $P_{h}^{v_{h}}(\boldsymbol{x})\in\mathcal{G}$ for all $\boldsymbol{x}\in\Omega$. This is not necessarily true for every $v_{h}\in V_{h}$. In the next result, we state the condition under which this fact can be achieved, and we will identify criteria that inform the assumptions made in Section 4.4.
Lemma 1.
Let $v_{h}\in V_{h}$. If $a-\underline{v}_{hi}\leq b-\overline{v}_{hi}$ for all node indices $i=1,\dots,N$, then $v^{+}_{h}(\boldsymbol{x})\in\mathcal{G}$ for all $\boldsymbol{x}\in\Omega$. .
3.1 The nonlinear finite element method
With the notations introduced in the last section, we introduce the finite element method studied in this work: Find $u_{h}\in V_{h}$ such that
$a_{h}(u^{+}_{h},v_{h})+s_{h}(u^{-}_{h},v_{h})=b_{h}(v_{h})\qquad\forall\,v_{h}\in V_{h}\,.$ (9)
Here, the bilinear form $a_{h}$ and linear form $b_{h}$ are defined as in (5), while the stabilization term for any $v_{h},w_{h}\in V_{h}$ reads
$s_{h}(w_{h},v_{h})=\alpha\sum_{i=1}^{N}(\epsilon h_{i}^{d-2}+\mu h_{i}^{d})w^{1}_{h}(\boldsymbol{x}_{i})v^{1}_{h}(\boldsymbol{x}_{i}).$ (10)
Here, $\alpha>0$ is a nondimensional stabilization parameter, and $h_{i}=\max\{h_{T}\mid T\in{\mathcal{T}_{h}}\text{ with }T\subset\omega_{i}\}$. The stabilization term $s_{h}$ induces the norm $\|\cdot\|_{s_{h}}$ on $V_{h}^{1}$ (and thus a seminorm in $V_{h}$).
The finite element method just presented is nonlinear since $u_{h}^{+}$ is a nonlinear transformation of $u_{h}$. So, it requires an appropriate linearization. In addition, as mentioned earlier, the presence of the term $h_{F}^{-\beta}$ in the stabilization can affect the condition number of the linear subproblems tremendously. To address these two issues, we now present a fixed-point iterative algorithm that enables us to both prove the existence of solutions and circumvent the potential ill-conditioning of the scheme. For this, we define the mapping
$\widetilde{T}\colon V^{0}_{h}\ni w^{0}_{h}\mapsto w^{1}_{h}\mapsto\tilde{w}^{0}_{h}\in V^{0}_{h}\,,$ (11a)
by the following algorithm:
Step 1: Compute $w^{1}_{h}\in V_{h}^{1}$ by solving the nonlinear scheme: $a_{h}([P^{w^{0}_{h}}_{h}w_{h}]^{1},v^{1}_{h})+s_{h}([Q^{w^{0}_{h}}_{h}w_{h}]^{1},v^{1}_{h})=(f,v^{1}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{0}_{h},v^{1}_{h})\qquad\forall v^{1}_{h}\in V^{1}_{h}.$ (11b) Step 2: Compute $\tilde{w}^{0}_{h}\in V^{0}_{h}$ as solution of $a_{h}(\tilde{w}^{0}_{h},v^{0}_{h})=(f,v^{0}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{1+}_{h},v^{0}_{h})-\underbrace{s_{h}(w^{1-}_{h},v^{0}_{h})}_{=0}\qquad\forall v^{0}_{h}\in V^{0}_{h}.$ (11c)
In the analysis presented in the next section, we will show that the operator $\widetilde{T}$ is well-defined. In addition, we make the following observations:
  1. 1.
    a fixed-point of $\widetilde{T}$ solves (9), which we can easily see by adding (11b) and (11c). So, in the next section we will show that $\widetilde{T}$ has a fixed point;
  2. 2.
    since the problem (11b) is posed over the space $V_{h}^{1}$ (which contains only continuous functions), the jump terms vanish. Thus, the condition number of the linear problems needed to solve (11b) is independent of $\beta$;
  3. 3.
    the left-hand side of (11c) can be simplified to
    $(\mu\tilde{w}_{h}^{0},v_{h}^{0})_{\mathcal{T}_{h}}+\sum_{F\in{\mathcal{F}_{h}}}\frac{\gamma(\epsilon+\mu h_{F}^{2})}{h_{F}^{\beta}}\int_{F}[\![\tilde{w}_{h}^{0}]\!]\cdot[\![v_{h}^{0}]\!]\,.$
    Since the mesh is assumed to be shape-regular, then the matrix associated with (11c) can be proven to have a condition number independent of $\beta$. We shall provide more details on this in Section 6.2.
4 Existence of a discrete solution
This section establishes the existence and stability of a solution to (9). To this end, we state our main result and prove it in the remainder of this section.
Theorem 1.
Let $\beta\geq 2$ and let us suppose that $\alpha,\gamma>0$ are sufficiently large. Then, there exists a solution $u_{h}=u_{h}^{1}+u_{h}^{0}\in V_{h}$ of (9). In addition, the piecewise constant part $u_{h}^{0}$ satisfies the following a priori bound
$\|[\![u^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]\,,$ (12)
where $C>0$ depends only on the shape-regularity of the mesh.
4.1 Preliminaries and auxiliary results
We exploit our fixed-point iteration (11a) to prove this result. Thus, we need to prove that (11b) and (11c) induce well-posed problems, which we perform in Lemmas 6 and 8, respectively. To this end, we need several auxiliary results.
Lemma 2 (Broken Poincaré inequality, [Bre03]).
For every $v^{0}_{h}\in V^{0}_{h}$ the following broken Poincaré inequality holds
$\|v^{0}_{h}\|_{0,\Omega}\leq\,C_{P}\left\{\sum_{F\in{\mathcal{F}_{h}}}\frac{1}{h_{F}}\|[\![v_{h}^{0}]\!]\|^{2}_{0,F}\right\}^{\frac{1}{2}}\,,$
where $C_{P}>0$ depends only on $\Omega$ and the regularity of the mesh.
Lemma 3 (Relation of bilinear forms, (3.8) in [BGPV24]).
There is a constant $C_{\textup{equiv}}$, which only depends on the mesh regularity, such that
$\left(\mu\|v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}+\epsilon\|\nabla v_{h}^{1}\|_{\mathcal{T}_{h}}^{2}\right)^{1/2}=\|v^{1}_{h}\|_{a_{h}}\leq\frac{C_{\textup{equiv}}}{\alpha}\|v^{1}_{h}\|_{s_{h}}\qquad\text{ for all }v^{1}_{h}\in V^{1}_{h}.$
Lemma 4.
Given $v^{0}_{h}\in V^{0}_{h}$, there are constants $C$ and $\tilde{c}$ (independent of $h$ and $v^{0}_{h}$) such that for any $T\in{\mathcal{T}_{h}}$
$\tilde{c}h_{T}^{2}|v^{1+}_{h}|^{2}_{1,T}\leq\|v^{1+}_{h}\|^{2}_{0,T},\quad\text{ and }\quad\|v^{1+}_{h}\|^{2}_{0,\Omega}\leq C\left[\max(|a|,|b|)^{2}|\Omega|+\|v^{0}_{h}\|^{2}_{0,\Omega}\right]\qquad\text{ for all }v^{1}_{h}\in V^{1}_{h}.$
Proof.
The first inequality follows from a standard local inverse inequality (cf. [EG21a, Lemma 12.1]). To show the second inequality, we start by using the mesh regularity and the result proven in [EG21a, Proposition 12.5], and get to
$\|v^{1+}_{h}\|^{2}_{0,\Omega}\leq\,C\sum_{i=1}^{N}h_{i}^{d}\,[v^{1+}_{h}(\boldsymbol{x}_{i})]^{2}\,.$ (13)
Next, the definition of $v^{1+}_{h}(\boldsymbol{x}_{i})$ implies that $|v^{1+}_{h}(\boldsymbol{x}_{i})|\leq\max(|a|,|b|)+\|v^{0}_{h}\|_{\infty,\omega_{i}}$. Using this fact, the inverse inequality, and the mesh regularity we get to
$\|v^{1+}_{h}\|^{2}_{0,\Omega}\leq\,C\sum_{i=1}^{N}[\max(|a|,|b|)+\|v^{0}_{h}\|_{\infty,\omega_{i}}]^{2}h_{i}^{d}\leq\,C\left[\max(|a|,|b|)^{2}|\Omega|+\|v^{0}_{h}\|^{2}_{0,\Omega}\right],$
which finishes the proof. ∎
In addition, we also recall the following trace-inverse inequality (see [EG21a, Lemma 12.10]): There exists $C>0$, depending only on the mesh regularity such that for any $T\in{\mathcal{T}_{h}}$ and a facet $F$ of $T$, and any $v_{h}\in V_{h}$ the following holds
$\|v_{h}\|_{0,F}\leq C\,h_{T}^{-1/2}\|v_{h}\|_{0,T}\,.$ (14)
The next two results are instrumental in the proof that the nonlinear problem from Step 1 (cf. (11b)) is well-posed. The first one is an extension of [BGPV24, Lem. 3.1] to the present (more involved) case, but the proof is identical to that of that result, and thus we omit it.
Lemma 5.
For given $w^{0}_{h}\in V^{0}_{h}$, we have the following relations for all $r^{1}_{h},v^{1}_{h}\in V^{1}_{h}$
$s_{h}([Q^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}-[Q^{w^{0}_{h}}_{h}v^{1}_{h}]^{1},[P^{w^{0}_{h}}_{h}r^{1}_{h}]^{1}-[P^{w^{0}_{h}}_{h}v^{1}_{h}]^{1})\geq 0.$
Lemma 6.
For given $w^{0}_{h}\in V^{0}_{h}$ let us define the operator $\tilde{T}_{1}\colon V^{1}_{h}\to[V^{1}_{h}]^{\ast}\simeq V^{1}_{h}$ via
$[\tilde{T}_{1}r^{1}_{h},v^{1}_{h}]=a_{h}([P^{w^{0}_{h}}_{h}r_{h}]^{1},v^{1}_{h})+s_{h}([Q^{w^{0}_{h}}_{h}r_{h}]^{1},v^{1}_{h})\qquad\text{ for all }v_{h}\in V^{1}_{h}.$ (15)
Then, $\tilde{T}_{1}$ is continuous. Moreover, if $\alpha\geq C_{\textup{equiv}}$, $\tilde{T}_{1}$ is strongly monotone in the sense that there is $C_{M}>0$ with
$[\tilde{T}_{1}v^{1}_{h}-\tilde{T}_{1}r^{1}_{h},v^{1}_{h}-r^{1}_{h}]\geq C_{M}\|v^{1}_{h}-r^{1}_{h}\|^{2}_{a_{h}}\qquad\text{ for all }v^{1}_{h},r^{1}_{h}\in V^{1}_{h}.$ (16)
Proof.
The proof follows exactly the same lines as that of [BGPV24, Th. 3.2] with [BGPV24, Lem. 3.1: (25)] replaced by Lemma 5. ∎
4.2 Well-posedness of the iteration defined in (11a)
Lemma 7 (Well-posedness of (11b)).
Under the assumptions of Lemma 6, equation (11b) uniquely defines $w^{1}_{h}\in V^{1}_{h}$ for a given $w^{0}_{h}$. Moreover, if $\hat{w}^{0}_{h},\bar{w}^{0}_{h}\in V^{0}_{h}$, and $\hat{w}^{1}_{h},\bar{w}^{1}_{h}\in V^{1}_{h}$ are their images under $\tilde{T}_{1}$, then the following Lipschitz continuity holds
$\|\hat{w}^{1}_{h}-\bar{w}^{1}_{h}\|_{a_{h}}\leq\tfrac{1}{C_{M}}\|\hat{w}^{0}_{h}-\bar{w}^{0}_{h}\|_{a_{h}}.$
Proof.
The existence and uniqueness of the solution of (11b) are a direct consequence of the strong monotonicity and continuity of $\tilde{T}_{1}$ (see, e.g., [RR06, Theorem 10.49]). To prove the Lipschitz continuity, let $\hat{w}_{h}^{0},\bar{w}_{h}^{0}\in V_{h}^{0}$ and $\hat{w}_{h}^{1},\bar{w}_{h}^{1}\in V_{h}^{1}$ be their images under $\tilde{T}_{1}$. Then, using (16) and the problem (11b) we get to
$C_{M}\|\hat{w}^{1}_{h}-\bar{w}^{1}_{h}\|_{a_{h}}^{2}\leq[\tilde{T}_{1}\hat{w}^{1}_{h},\hat{w}^{1}_{h}-\bar{w}^{1}_{h}]-[\tilde{T}_{1}\bar{w}^{1}_{h},\hat{w}^{1}_{h}-\bar{w}^{1}_{h}]=a_{h}(\bar{w}^{0}_{h}-\hat{w}^{0}_{h},\hat{w}^{1}_{h}-\bar{w}^{1}_{h}),$
and the proof is finished using the continuity of $a_{h}(\cdot,\cdot)$. ∎
Once the proof that Step 1 is well-defined, we prove that the second step in the definition of $\widetilde{T}$ is also well-posed, thus proving that $\widetilde{T}$ is well-defined.
Lemma 8 (Well-posedness of (11c)).
For any given $w^{1}_{h}\in V^{1}_{h}$, there exists a unique $\tilde{w}^{0}_{h}\in V^{0}_{h}$ solution of (11c). Moreover, if $\hat{w}^{1+}_{h},\bar{w}^{1+}_{h}\in V^{1}_{h}$ and $\hat{w}^{0}_{h},\bar{w}^{0}_{h}\in V^{0}_{h}$ are the solutions of (11c) with $\hat{w}^{1+}_{h}$ and $\bar{w}^{1+}_{h}$ as right-hand sides, respectively, the following Lipschitz continuity holds
$\|\hat{w}^{0}_{h}-\bar{w}^{0}_{h}\|_{a_{h}}\leq\|\hat{w}^{1+}_{h}-\bar{w}^{1+}_{h}\|_{a_{h}}.$
Proof.
For $w_{h}^{0},v_{h}^{0}\in V_{h}^{0}$ the bilinear form on the left-hand side of (11c) reduces to
$a_{h}(w^{0}_{h},v^{0}_{h})=\mu(\tilde{w}^{0}_{h},v^{0}_{h})_{\mathcal{T}_{h}}+\gamma\left\langle\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![\tilde{w}^{0}_{h}]\!],[\![v^{0}_{h}]\!]\right\rangle_{\mathcal{F}_{h}}\qquad\text{ for all}w^{0}_{h},v^{0}_{h}\in V^{0}_{h}.$ (17)
This form is continuous and $\|[\![\cdot]\!]\|_{\mathcal{F}_{h}}$-elliptic with ellipticity constant $\gamma h^{1-\beta}$. Thus, the problem in (11c) is well-posed. To prove the Lipschitz continuity, we consider (11c) with the two different right-hand sides, subtract the equations, and get to
$a_{h}(\hat{w}^{0}_{h}-\bar{w}^{0}_{h},v_{h}^{0})=a_{h}(\hat{w}^{1+}_{h}-\bar{w}^{1+}_{h},v_{h}^{0}),$
for every $v_{h}^{0}\in V_{h}^{0}$. Taking $v_{h}^{0}=\hat{w}^{0}_{h}-\bar{w}^{0}_{h}$ and using Cauchy-Schwarz’s inequality on the right-hand side above leads to the result. ∎
4.3 Proof of Theorem 1
We start proving the following a priori stability result for the solution of (11c).
Lemma 9.
Let $\beta\geq 2$, $w^{1}_{h}\in V^{1}_{h}$ be given, and let $\tilde{w}^{0}_{h}\in V^{0}_{h}$ be the solution of (11c) with $w^{1}_{h}\in V^{1}_{h}$ on the right-hand side. Then, there exists a constant $C>0$, depending only on the mesh regularity, such that
$\|[\![\tilde{w}^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\|w^{1+}_{h}\|^{2}_{0,\Omega}\right].$
Proof.
Let $v^{0}_{h}=\tilde{w}^{0}_{h}$ in (11c) and $\tilde{\gamma}=\tfrac{\gamma}{h^{\beta-1}}$. Then, using (17), the ellipticity of $a_{h}(\cdot,\cdot)$ in $V_{h}^{0}$, (11c), Cauchy-Schwarz’s inequality, the local trace-inverse inequality (14), the broken Poincaré inequality from Lemma 2, and a global inverse inequality, we get to
$\|\sqrt{\mu}\tilde{w}^{0}_{h}\|^{2}_{0,\Omega}+\tilde{\gamma}\|[\![\tilde{w}^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq a_{h}(\tilde{w}^{0}_{h},\tilde{w}^{0}_{h})$
$=(f,\tilde{w}^{0}_{h})_{\mathcal{T}_{h}}-a_{h}(w^{1+}_{h},\tilde{w}^{0}_{h})$
$=(f,\tilde{w}^{0}_{h})_{\mathcal{T}_{h}}-(\mu w^{1+}_{h},\tilde{w}^{0}_{h})_{\mathcal{T}_{h}}-\langle\{\!\!\{\varepsilon\nabla w_{h}^{1+}\}\!\!\},[\![\tilde{w}^{0}_{h}]\!]\rangle_{\mathcal{F}_{h}}$
$\leq C_{P}\mu^{-\frac{1}{2}}h^{-1}\|f\|_{0,\Omega}\|[\![\tilde{w}^{0}_{h}]\!]\|_{\mathcal{F}_{h}}+C_{P}\sqrt{\mu}h^{-1}\|w^{1+}_{h}\|_{0,\Omega}\|[\![\tilde{w}^{0}_{h}]\!]\|_{\mathcal{F}_{h}}+C\sqrt{\epsilon}\|\nabla w_{h}^{1+}\|_{0,\Omega}\|[\![\tilde{w}^{0}_{h}]\!]\|_{\mathcal{F}_{h}}$
$\leq C\Big{(}C_{P}\mu^{-\frac{1}{2}}h^{-1}\|f\|_{0,\Omega}+C_{P}\sqrt{\mu}h^{-1}\|w^{1+}_{h}\|_{0,\Omega}+\sqrt{\epsilon}h^{-1}\|w_{h}^{1+}\|_{0,\Omega}\Big{)}\|[\![\tilde{w}^{0}_{h}]\!]\|_{\mathcal{F}_{h}}\,.$
The proof then follows from rearranging terms. ∎
Lemma 10.
Let us assume that $\beta\geq 2$. Then, if $\gamma>0$ is sufficiently large, $T\colon B_{1}(0)\to B_{1}(0)$, where $B_{1}(0)\subset V^{0}_{h}$ is the closed unit ball in $(V_{h}^{0},\|[\![\cdot]\!]\|_{\mathcal{F}_{h}})$.
Proof.
Using Lemma 4, Lemma 9, and the broken Poincaré inequality from Lemma 2 we get to
$\|[\![\tilde{w}^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\|w^{1+}_{h}\|^{2}_{0,\Omega}\right]$
$\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\big{(}\max(|a|,|b|)^{2}|\Omega|+\|w^{0}_{h}\|^{2}_{0,\Omega}\big{)}\right]$
$\leq\,C\,\frac{h^{2\beta-4}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|+C_{P}^{2}\frac{C_{p}^{2}\mu+\epsilon}{\epsilon+\mu h^{2}}\|[\![w^{0}_{h}]\!]\|^{2}_{{\mathcal{F}_{h}}}\right]\,.$ (18)
The result is then proved by noticing that $\gamma$ can always be chosen large enough so that the last term in the right-hand side above can be hidden in the left-hand side, and the resulting right-hand side is smaller than $1$. ∎
The proof of Theorem  1 appears as a corollary of the above results.
Proof of Theorem 1.
Since the EG finite element space is finite-dimensional, the operator $\tilde{T}$ is continuous and maps the closed unit ball (a convex compact set) onto itself. Thus, Brouwer’s fixed-point Theorem (see, e.g. [RR06, Theorem 10.41]) states that $\tilde{T}$ has at least one fixed point in $B_{1}(0)$, which is the desired result. ∎
4.4 Worst-case criterion for positivity preservation
Corollary 1 implies that for each iterate $w_{h}\in V_{h}$ defined via (11a), we have $w^{+}_{h}\in\mathcal{G}$ if $\|w^{0}_{h}\|_{0,\infty,\Omega}<\frac{b-a}{2}$. This section underlines that this criterion can always easily be satisfied if either $\beta$ is large enough, $\gamma$ is large enough, or $h$ is small enough. To this end, we start observing that using the inverse inequality, we get
$\|\tilde{w}^{0}_{h}\|_{0,\infty,\Omega}\leq C\,h^{-d/2}\|\tilde{w}^{0}_{h}\|_{0,\Omega},$
which can be combined with Lemma 2 and (4.3) to obtain
$\|\tilde{w}^{0}_{h}\|_{0,\infty,\Omega}\leq\,C\frac{h^{\beta-2-d/2}}{\gamma}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\big{(}\max(|a|,|b|)^{2}|\Omega|+\|w_{h}^{0}\|_{0,\Omega}^{2}\big{)}\right]^{\frac{1}{2}}\,.$
So, assuming that $\beta\geq 2+d/2$, or that $\gamma$ is large enough, for $w_{h}^{0}\in V_{h}^{0}$ satisfying $\|w^{0}_{h}\|^{2}_{0,\Omega}\leq 1$, we conclude that $\|\tilde{w}^{0}_{h}\|_{0,\infty,\Omega}\leq(b-a)/2$, thus guaranteeing that the next iterate $\tilde{w}^{+}_{h}$ belongs to $\mathcal{G}$.
5 Convergence order estimates
In this section, we prove optimal order error estimates for the present bound-preserving EG method. The order of convergence appears as a result of the over-penalization of the jump terms. More precisely, the stability bound (12) proven in Theorem 1 is an error estimate in its own right, providing optimal convergence for $u_{h}^{0}$ if ${\beta\geq 4}$, since in such a case $u^{0}_{h}\to 0$ is sufficiently fast. Thus, the approximation properties of the current EG method stem from those of the continuous finite element subspace, while the piecewise constant enrichment is responsible for the local mass conservation. A similar interpretation of the best approximation properties can be used to argue for the convergence of DG methods, which converge optimally since the jump terms in the solutions converge to zero sufficiently quickly (the continuous solution does not have discontinuities).
We start by stating the following estimate for $u_{h}^{0}$. It is important to notice that this result is, in fact, a rewriting of Theorem 1.
Corollary 1.
Under the assumptions of Theorem 1, Let $u_{h}^{0}$ be the piecewise constant part of a solution $u_{h}{}$ of (9). Then, the following error estimate holds
$\|[\![u^{0}_{h}]\!]\|^{2}_{\mathcal{F}_{h}}\leq\,C\,\frac{h^{{2\beta-4}}}{\gamma^{2}}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]\,.$ (19)
Remark 3.
Thanks to the broken Poincaré inequality from Lemma 2, we also have that $\|u_{h}^{0}\|_{0,\Omega}\leq C\,h^{\beta-{3/2}}$.
We now state the main convergence result of this work.
Theorem 2.
Let us assume that $u\in H^{2}(\Omega)$ solves (2), $\beta\geq 4$, and that the assumptions of Theorem 1 hold. Then, the following error estimate holds
$\Big{(}\mu\|u-u^{+}_{h}\|_{0,\mathcal{T}_{h}}^{2}+\epsilon\|\nabla(u-u_{h}^{+})\|_{0,\mathcal{T}_{h}}^{2}+\|[\![u_{h}^{0}]\!]\|_{{\mathcal{F}_{h}}}^{2}\Big{)}^{1/2}\leq C\,h(\sqrt{\mu}h+\sqrt{\epsilon})\,\Big{(}|u|_{2,\Omega}+\sqrt{\tilde{C}(f,a,b,\Omega)}\Big{)}\,,$
where $\tilde{C}(f,a,b,\Omega)=\gamma^{-2}\left[C_{P}^{2}\mu^{-1}\|f\|^{2}_{0,\Omega}+(C_{P}^{2}\mu+\epsilon)\max(|a|,|b|)^{2}|\Omega|\right]$, and the constant $C$ does not depend on the mesh size, nor the physical coefficients of the problem.
The first step in the proof of the error is the following result that states that the constrained part of the discrete solution $u_{h}^{+}$ satisfies a variational inequality.
Lemma 11.
Let $u_{h}\in V_{h}$ solve (9) and define the closed convex set
$V^{+}_{h}=\{v_{h}\in V_{h}\colon v_{h}=P^{u^{0}_{h}}_{h}w_{h}\text{ for some }w_{h}\in V_{h}\}.$ (20)
Then, the function $u^{+}_{h}$ satisfies
$a_{h}(u_{h}^{+},v_{h}-u^{+}_{h})\geq b(v_{h}-u_{h}^{+})\qquad\forall v_{h}\in V^{+}_{h}.$
Proof.
As in the proof of [BGPV24, Th. 3.5], we can derive that
$a_{h}(u_{h}^{+},v_{h}-u^{+}_{h})+s(u^{1-}_{h},v^{1}_{h}-u^{1+}_{h})=b(v_{h}-u_{h}^{+})\qquad\forall v_{h}\in V_{h}^{+}.$
Using that $P^{u^{0}_{h}}_{h}v_{h}=v_{h}$ and $Q^{u^{0}_{h}}_{h}v_{h}=0$ for $v_{h}\in V_{h}^{+}$ we exploit Lemma 5 to deduce that
$s(u^{1-}_{h},v^{1}_{h}-u^{1+}_{h})=s([Q^{u^{0}_{h}}_{h}u_{h}]^{1}-[Q^{u^{0}_{h}}_{h}v_{h}]^{1},[P^{u^{0}_{h}}_{h}v_{h}]^{1}-[P^{u^{0}_{h}}_{h}u_{h}]^{1})\leq 0\qquad\forall v_{h}\in V^{+}_{h},$
which implies the result. ∎
To exploit the variational inequality proven above, we must choose an appropriate test function. To achieve this, we introduce the Lagrange interpolation operator $\mathcal{I}_{h}:C^{0}(\bar{\Omega})\to V_{h}^{1}$ defined in [EG21a, Chapter 11], and define the test function
$\tilde{v}_{h}:=P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u).$ (21)
Then, the following result holds.
Lemma 12.
The following error estimate holds
$\mu\|u-u_{h}^{+}\|_{0,\mathcal{T}_{h}}^{2}+\epsilon\|\nabla(u-u_{h}^{+})\|_{0,\mathcal{T}_{h}}^{2}\leq C\Big{(}\epsilon\,\|\nabla(u-\tilde{v}_{h})\|_{0,\mathcal{T}_{h}}^{2}+\mu\|u-\tilde{v}_{h}\|_{0,\mathcal{T}_{h}}^{2}+\|[\![u_{h}^{0}]\!]\|_{{\mathcal{F}_{h}}}^{2}\Big{)}\,,$
where $C>0$ depends only on the mesh regularity.
Proof.
Thanks to the regularity assumption on $u$, the EG method used in this work is consistent, and thus
$a_{h}(u,v_{h}-u^{+}_{h})=b(v_{h}-u^{+}_{h})\,,$
for every $v_{h}\in V_{h}$. So, for every $v_{h}\in V_{h}^{+}$, where we remind that $V_{h}^{+}$ is given by (20), the following inequality holds
$a_{h}(u_{h}^{+}-u,v_{h}-u^{+}_{h})\geq 0\,.$
Next, considering $\tilde{v}_{h}$ defined in (21), using the ellipticity of the bilinear form $a_{h}(\cdot,\cdot)$, the Cauchy-Schwarz inequality, the local trace-inverse result (14), the fact that $\tilde{v}_{h}-u_{h}^{+}$ is continuous, and that $[\![\tilde{v}_{h}-u]\!]={[\![P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u)]\!]=}[\![u_{h}^{0}]\!]$ we get
$\|\tilde{v}_{h}-u_{h}^{+}\|_{a_{h}}^{2}$ $=a_{h}(\tilde{v}_{h}-u_{h}^{+},\tilde{v}_{h}-u_{h}^{+})$
$=a_{h}(\tilde{v}_{h}-u,\tilde{v}_{h}-u_{h}^{+})+\underbrace{a_{h}(u-u_{h}^{+},\tilde{v}_{h}-u_{h}^{+})}_{\leq 0}$
$\leq a_{h}(\tilde{v}_{h}-u,\tilde{v}_{h}-u_{h}^{+})$
$=\epsilon(\nabla(\tilde{v}_{h}-u),\nabla(\tilde{v}_{h}-u_{h}^{+}))_{\mathcal{T}_{h}}+\mu(\tilde{v}_{h}-u,\tilde{v}_{h}-u_{h}^{+})_{\mathcal{T}_{h}}-\langle\epsilon\{\!\!\{\nabla(\tilde{v}_{h}-u_{h}^{+})\}\!\!\},[\![\tilde{v}_{h}-u]\!]\rangle_{\mathcal{F}_{h}}$
$\leq C\Big{(}\epsilon\|\nabla(u-\tilde{v}_{h})\|_{0,\mathcal{T}_{h}}^{2}+\mu\|u-\tilde{v}_{h}\|_{0,\mathcal{T}_{h}}^{2}+\|[\![u_{h}^{0}]\!]\|_{{\mathcal{F}_{h}}}^{2}\Big{)}^{\frac{1}{2}}\|\tilde{v}_{h}-u_{h}^{+}\|_{a_{h}}\,,$
which, after using the triangle inequality, proves the result. ∎
To prove the error, it only remains to bound the difference $u-\tilde{v}_{h}=u-P_{h}^{u_{h}^{0}}(\mathcal{I}_{h}u)$. We remark that $u_{h}^{0}$ has already been bounded in Corollary 1, so it only remains to bound the difference $u-{[P_{h}^{u_{h}^{0}}(\mathcal{I}_{h}u)]^{1}}$. The following result states that bound.
Lemma 13.
Let $w_{h}{=w_{h}^{1}+w_{h}^{0}}\in V_{h}$ be arbitrary. Then, there exists a constant $C>0$ depending only on the mesh regularity such that
${\|w^{1}_{h}-[P_{h}^{w_{h}^{0}}(w_{h})]^{1}\|_{0,\Omega}=}\|w^{1}_{h}-w^{1+}_{h}\|_{0,\Omega}$ $\leq C\,\|w_{h}^{0}\|_{0,\Omega},$
${\|\nabla w^{1}_{h}-\nabla[P_{h}^{w_{h}^{0}}(w_{h})]^{1}\|_{0,\Omega}=}\|\nabla w^{1}_{h}-\nabla w^{1+}_{h}\|_{\mathcal{T}_{h}}$ $\leq C\,h^{-1}\|w_{h}^{0}\|_{0,\Omega}.$
Proof.
Let $T\in{\mathcal{T}_{h}}$. Then, using (13) and the regularity of the mesh family, we get
$\|w^{1}_{h}-w^{1+}_{h}\|^{2}_{0,T}\leq C\,h^{d}\max_{\boldsymbol{x}_{i}\in T}(w^{1}_{h}(\boldsymbol{x}_{i})-w^{1+}_{h}(\boldsymbol{x}_{i}))^{2}\leq Ch^{d}\max_{\hat{}T\in\omega_{T}}\|w_{h}^{0}\|^{2}_{0,\infty,\omega_{T}}\leq C\,\|w_{h}^{0}\|^{2}_{0,\omega_{T}}.$
Adding over the elements in the mesh yields the first inequality. The second one follows using the inverse inequality. ∎
Gathering all these preliminary results, we can finally prove Theorem 2.
Proof of Theorem 2.
Classical approximation properties of $\mathcal{I}_{h}$ (see, e.g., [EG21a]) yield
$\sqrt{\mu}\|u-\mathcal{I}_{h}u\|_{0,\Omega}+\sqrt{\epsilon}\|\nabla(u-\mathcal{I}_{h}u)\|_{0,\Omega}\leq C\,(\sqrt{\mu}h+\sqrt{\epsilon})h|u|_{2,\Omega}\,.$
In addition, using Lemma 13, the approximation properties of $\mathcal{I}_{h}$ and Corollary 1 we arrive at
$\|u-\tilde{v}_{h}\|_{0,\Omega}$ $\leq\|u-\mathcal{I}_{h}u\|_{0,\Omega}+\|\mathcal{I}_{h}u-{[P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u)]^{1}}\|_{0,\Omega}+\|u_{h}^{0}\|_{0,\Omega}$
$\leq Ch^{2}|u|_{2,\Omega}+C\|u_{h}^{0}\|_{0,\Omega}$
$\leq Ch^{2}|u|_{2,\Omega}+Ch^{\beta-2}\sqrt{\tilde{C}(f,a,b,\Omega)}\,,$
$\|\nabla(u-\tilde{v}_{h})\|_{0,\mathcal{T}_{h}}$ $\leq\|\nabla(u-\mathcal{I}_{h}u)\|_{0,\mathcal{T}_{h}}+\|\nabla(\mathcal{I}_{h}u-{[P^{u^{0}_{h}}_{h}(\mathcal{I}_{h}u)]^{1}})\|_{0,\mathcal{T}_{h}}$
$\leq Ch|u|_{2,\Omega}+Ch^{-1}\|u_{h}^{0}\|_{0,\Omega}$
$\leq Ch|u|_{2,\Omega}+Ch^{\beta-3}\sqrt{\tilde{C}(f,a,b,\Omega)}\,.$
Gathering the last inequalities and using that $\beta\geq 4$ proves the error estimate. ∎
6 Numerical examples
This section presents various numerical examples to validate our theoretical findings and discuss implementation aspects. All computations were done using the finite element library Netgen/NGSovle, see [Sch14], www.ngsolve.org.
6.1 An iterative algorithm
To solve the nonlinear problem (9), we follow a similar strategy as was presented in [BGPV24], i.e., we consider using a Richardson-like iterative approximation. However, since an iteration of the (linearized) fully coupled problem defined on $V_{h}$ might lead to very ill-conditioned system matrices, see Section 6.2 below, we define a nested iterative scheme that is motivated by the fix point iteration (11a) introduced in Section 3. Each step in the iterative algorithm includes solving equations (11b) and (11c). While (11c) is linear and can be solved with classical tools (e.g., a direct solver), we introduce another iterative (sub) scheme to solve the non-linear problem (11b).
To track the nested iterative algorithm, we use the following notation. The outer loop, resembling the iteration of the fixed point iteration (11a), is denoted by the index $m$. The inner loop, resembling the iteration of the linearized problem (11b), is denoted by the index $n$. Correspondingly, we denote use $u_{h}^{1}\big{|}^{n}_{m}$ to indicate the linear part of the solution of the linearized problem at the $n$-th inner iteration of the $m$-th outer iteration. Since $u_{h}\big{|}_{m}$ is fixed in the inner loop, we omit the index $m$ for the constants in the following. Now consider a tolerance $tol_{m}$ for the outer loop and a tolerance $tol_{n}$ for the inner loop. The algorithm is then defined as follows:
Given a starting guess $u_{h}^{0}\big{|}_{0}\in V_{h}^{0}$ we solve for all $m=0,1,...$ the problem for $u_{h}^{0}\big{|}_{m+1}$: Step 1: Solve the following iteration to find $u_{h}^{1}\big{|}_{m+1}$: Given the starting guess $u_{h}^{1}\big{|}_{m}^{0}=u_{h}^{1}\big{|}_{m}$ and a damping parameter $\omega\in(0,1]$, for each $n=0,...$ we find $u_{h}^{1}\big{|}^{n+1}_{m}\in V_{h}^{1}$ such that for all $v_{h}^{1}\in V_{h}^{1}$ there holds $a_{h}(u_{h}^{1}\big{|}^{n+1}_{m},v_{h}^{1})=$ $a_{h}(u_{h}^{1}\big{|}_{m}^{n},v_{h}^{1})$ $+\omega\Big{(}(f,v_{h}^{1})-a_{h}([P_{h}^{u_{h}^{0}\big{|}_{m}}u_{h}^{1}\big{|}_{m}^{n}]^{1},v_{h}^{1})-s_{h}([Q_{h}^{u_{h}^{0}\big{|}_{m}}u_{h}^{1}|^{n}_{m}]^{1},v_{h}^{1})\Big{)},$ until $\|u_{h}^{1}\big{|}^{n+1}_{m}-u_{h}^{1}\big{|}^{n}_{m}\|_{0}\leq tol_{n}$. Let $N$ denote the number of iterations needed to reach the tolerance $tol_{n}$. We then set $u_{h}^{1}\big{|}_{m+1}=u_{h}^{1}\big{|}^{N}_{m}$. Step 2: To find $u_{h}^{0}\big{|}_{m+1}$, solve the problem: $a_{h}(u_{h}^{0}\big{|}_{m+1},v_{h}^{0})=(f,v_{h}^{0})-a_{h}(u_{h}^{1}\big{|}_{m},v_{h}^{0}),\quad\forall v_{h}^{0}\in V_{h}^{0}.$ Terminate the (outer) iteration when $\|u_{h}^{0}\big{|}_{m+1}-u_{h}^{0}\big{|}_{m}\|_{0}\leq tol_{m}$.
We chose $\omega=1$ in our numerical examples and tested different values for the stopping criteria. Surprisingly, we found that the tolerance of the inner loop is not crucial for the algorithm’s convergence. We observed that the algorithm converges for a wide range of tolerances; see Section 6.3 for more details. As a starting guess for the outer loop, we used $u_{h}^{0}\big{|}_{0}=[u_{h}^{init}]^{0}$, i.e., the constant part of $u_{h}^{init}$ given as the solution of the standard EG problem without modification, i.e.
$a_{h}(u_{h}^{init},v_{h})=(f,v_{h}),\quad\forall v_{h}\in V_{h}.$
We similarly used $u_{h}^{1}\big{|}_{0}=[u_{h}^{init}]^{1}$ for the inner loop.
6.2 The condition number of the linearized problems
Let $\mathbb{A}$ denote the finite element matrix associated to the bilinear form $a_{h}(\cdot,\cdot)$, that is, the matrix that needs to be inverted when solving a linearized problem of (9). In the discussion that follows, we can neglect the stabilization term $s_{h}(\cdot,\cdot)$, as it only appears on the right-hand side of the inner iterations in Step 1 of the scheme presented in the last section. Further let $a_{h}^{i}(\cdot,\cdot)$ with $i\in\{0,1\}$ denote the restriction of the bilinear form $a_{h}$ on $V_{h}^{i}$, and let $\mathbb{A}^{i}$ denote the corresponding matrix. In the following, we discuss the condition numbers we expect for each system matrix, considering varying choices of $\beta$. By construction, the matrix $\mathbb{A}^{1}$ is the standard finite element matrix of linear Lagrange finite elements and is independent of $\beta$. The condition $\kappa(\mathbb{A}^{1})$ is then known to scale like $\mathcal{O}(h^{-2})$ which follows with standard techniques, see e.g. [EG21b, Chap. 28].
In the forthcoming discussion, we will write $a\lesssim b$ to denote that $a\leq C\,b$, where $C>0$ is a constant independent of $h$ (but that might depend on the physical parameters $\mu$ and $\epsilon$), and $a\sim b$ if $a\lesssim b$ and $b\lesssim a$. To discuss the condition number of $\mathbb{A}^{0}$ we introduce the bilinear forms
$m(w_{h}^{0},v_{h}^{0})=\int_{\Omega}w_{h}^{0}v_{h}^{0}\;\textup{d}\boldsymbol{x},\quad\text{and}\quad j(w_{h}^{0},v_{h}^{0})=\langle[\![w^{0}_{h}]\!],[\![v^{0}_{h}]\!]\rangle,$
for $w_{h}^{0},v_{h}^{0}\in V_{h}^{0}$, and denote by $\mathbb{M}$ and $\mathbb{J}$ the corresponding system matrices. Then, assuming for simplicity $\mu=\gamma=1$, we have that $\mathbb{A}^{0}=\mathbb{M}+h^{-\beta}\mathbb{J}$ (with the obvious abuse of notation regarding the term containing the negative power of $h$). In the first step, we only consider the jump matrix $\mathbb{J}$. Let $u_{h}^{0}\in V_{h}^{0}$ be arbitrary, and let $\underline{u}_{0}$ denote the corresponding finite element coefficient vector.
Using the broken Poincaré inequality from Lemma 2 and the inverse trace inequality (14) we obtain
$\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}=\|u_{h}^{0}\|^{2}_{0,\Omega}\lesssim\underbrace{h^{-1}\langle[\![u^{0}_{h}]\!],[\![u^{0}_{h}]\!]\rangle}_{h^{-1}\underline{u}_{0}^{T}\mathbb{J}\underline{u}_{0}}\lesssim h^{-2}\|u_{h}^{0}\|^{2}_{0,\Omega}=h^{-2}\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0},$
that is, after multiplying by $h$,
$h\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\lesssim\underline{u}_{0}^{T}\mathbb{J}\underline{u}_{0}\lesssim h^{-1}\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\,,$ (22)
and using that $\kappa(\mathbb{M})\sim\mathcal{O}(1)$ (see, e.g., [EG21b, Proposition 28.6]) we conclude that $\kappa(\mathbb{J})\sim\mathcal{O}(h^{-2})$ (and thus is independent of the value of $\beta$). This further implies
$(1+h^{-\beta+1})\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0}\lesssim\underbrace{\underline{u}_{0}^{T}(\mathbb{M}+h^{-\beta}\mathbb{J})\underline{u}_{0}}_{\sim\underline{u}_{0}^{T}\mathbb{A}^{0}\underline{u}_{0}}\lesssim(1+h^{-(\beta+1)})\underline{u}_{0}^{T}\mathbb{M}\underline{u}_{0},$
from which we also conclude $\kappa(\mathbb{J})\sim\mathcal{O}\Big{(}(1+h^{-(\beta+1)})/(1+h^{-\beta+1})\Big{)}$ which is bounded by $Ch^{-2}$ independently of the choice of $\beta$.
Numerical investigation:
To validate the findings presented above, we computed the condition numbers for a structured triangulation of the domain $\Omega=(0,1)^{2}$ with various mesh sizes. The results, summarized in Table 1, demonstrate that $\kappa(\mathbb{A}),\kappa(\mathbb{A}^{0})$ and $\kappa(\mathbb{A}^{1})$ scale at the anticipated rates, consistent with the theoretical predictions discussed earlier.
$\beta=1$
$h$ $\kappa(\mathbb{A})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{1})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{0})^{-1}$ $eoc$
$0.5$ $1.52\cdot 10^{-3}$ (--- -) $2.03\cdot 10^{-2}$ (--- -) $1.28\cdot 10^{-1}$ (--- -)
$0.25$ $5.14\cdot 10^{-4}$ ($1.57$) $5.60\cdot 10^{-3}$ ($1.86$) $4.12\cdot 10^{-2}$ ($1.63$)
$0.125$ $1.55\cdot 10^{-4}$ ($1.73$) $1.59\cdot 10^{-3}$ ($1.82$) $1.20\cdot 10^{-2}$ ($1.78$)
$0.0625$ $4.31\cdot 10^{-5}$ ($1.85$) $4.36\cdot 10^{-4}$ ($1.87$) $3.25\cdot 10^{-3}$ ($1.88$)
$0.0313$ $1.14\cdot 10^{-5}$ ($1.92$) $1.15\cdot 10^{-4}$ ($1.92$) $8.49\cdot 10^{-4}$ ($1.94$)
$\beta=2$
$h$ $\kappa(\mathbb{A})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{1})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{0})^{-1}$ $eoc$
$0.5$ $6.23\cdot 10^{-4}$ (--- -) $2.03\cdot 10^{-2}$ (--- -) $1.06\cdot 10^{-1}$ (--- -)
$0.25$ $1.06\cdot 10^{-4}$ ($2.55$) $5.60\cdot 10^{-3}$ ($1.86$) $3.53\cdot 10^{-2}$ ($1.59$)
$0.125$ $1.61\cdot 10^{-5}$ ($2.72$) $1.59\cdot 10^{-3}$ ($1.82$) $1.04\cdot 10^{-2}$ ($1.76$)
$0.0625$ $2.24\cdot 10^{-6}$ ($2.84$) $4.36\cdot 10^{-4}$ ($1.87$) $2.84\cdot 10^{-3}$ ($1.87$)
$0.0313$ $2.97\cdot 10^{-7}$ ($2.92$) $1.15\cdot 10^{-4}$ ($1.92$) $7.43\cdot 10^{-4}$ ($1.93$)
$\beta=4$
$h$ $\kappa(\mathbb{A})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{1})^{-1}$ $eoc$ $\kappa(\mathbb{A}^{0})^{-1}$ $eoc$
$0.5$ $9.51\cdot 10^{-5}$ (--- -) $2.03\cdot 10^{-2}$ (--- -) $6.69\cdot 10^{-2}$ (--- -)
$0.25$ $4.15\cdot 10^{-6}$ ($4.52$) $5.60\cdot 10^{-3}$ ($1.86$) $2.35\cdot 10^{-2}$ ($1.51$)
$0.125$ $1.58\cdot 10^{-7}$ ($4.71$) $1.59\cdot 10^{-3}$ ($1.82$) $7.09\cdot 10^{-3}$ ($1.73$)
$0.0625$ $5.52\cdot 10^{-9}$ ($4.84$) $4.36\cdot 10^{-4}$ ($1.87$) $1.96\cdot 10^{-3}$ ($1.85$)
$0.0313$ $1.83\cdot 10^{-10}$ ($4.92$) $1.15\cdot 10^{-4}$ ($1.92$) $5.16\cdot 10^{-4}$ ($1.92$)
Table 1: Condition numbers $\kappa(\mathbb{A}),\kappa(\mathbb{A}^{0})$ and $\kappa(\mathbb{A}^{1})$ obtained with a structured triangulation of the domain $\Omega=(0,1)^{2}$ with varying mesh sizes $h$.
6.3 Smooth solution
Let $\Omega=(-1,1)\times(0,1)$ and the parameters be set to $\epsilon=10^{-5}$ and $\mu=1$. Further let $f$ be defined such that the exact solution of (2) is given by
$u(x,y)=\sin(\pi(x+1)/2)\cdot\sin(\pi y).$
i.e., we have $a=0$ and $b=1$. In the following, we investigate the convergence of the method on a sequence of (nested) unstructured triangulations $\mathcal{T}_{h}$. As suggested by the theory, we choose $\beta=4$ (for the first two test cases) and the tolerance of the outer loop is $tol_{m}=10^{-12}$. Further, following [BGPV24], we choose the damping parameter $\omega=0.5$ and set $\alpha=1$ in (10).
Convergence of the method and choice of $tol_{n}$:
In Table 2, we present the results for the convergence of the method for different tolerances of the inner loop and the stabilization factor $\gamma=10$. We observe that the method converges for a wide range of tolerances of the inner loop, and the number of (outer) iterations $it$ is only mildly influenced by the choice of $tol_{n}$. Further, we see that the estimated order of convergence $eoc$ of the errors is optimal, i.e., we observe a quadratic rate for the error $\|u-u_{h}^{+}\|_{0,\Omega}$, and a linear rate for $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$. Note, that the $H^{1}$-error only considers the linear part in $V_{h}^{1}$ but not the constant part in $V_{h}^{0}$, i.e. $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}=\|\nabla(u-(u_{h}^{+})^{1})\|_{0,\Omega}$.
$tol_{n}=10^{-3}$
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ $eoc$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$$\#its$
$496$ $5.33\cdot 10^{-3}$ (--- -) $3.44\cdot 10^{-1}$ (--- -) $9$
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) $8$
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) $6$
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) $2$
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) $1$
$tol_{n}=10^{-6}$
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ $eoc$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$$\#its$
$496$ $5.33\cdot 10^{-3}$ (--- -) $3.44\cdot 10^{-1}$ (--- -) $5$
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) $5$
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) $4$
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) $2$
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) $1$
$tol_{n}=10^{-9}$
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ $eoc$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$$\#its$
$496$ $5.33\cdot 10^{-3}$ (--- -) $3.44\cdot 10^{-1}$ (--- -) $3$
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) $2$
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) $2$
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) $2$
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) $1$
Table 2: Error convergence and number of iterations for the example of Section 6.3 with $tol_{m}=10^{-12},\gamma=10,\beta=4$ and varying tolerances $tol_{n}$.
The choice of $\gamma$:
In Figure 1 we present the number of iterations for different stabilization factors $\gamma$ and a tolerance of the inner loop set to $tol_{n}=10^{-9}$. In the left plot, the numbers are given for $\epsilon=10^{-5}$ and on the right for $\epsilon=1$. We observe that the number of iterations is barely influenced by the choice of $\gamma$, even for a small $\epsilon$. To motivate this behavior, we recall the stabilizing term in the bilinear form $a_{h}(\cdot,\cdot)$ given by $\langle\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\rangle$. We see that even in the case of small $\epsilon$ and a coarse mesh, since $\mu=1$ and $\beta\geq 4$, we get a relatively big penalization of the jumps of the solution even for a moderate $\gamma$. This would not be the case (particularly for vanishing $\epsilon$) if we use the stabilization $\gamma\frac{\epsilon}{h_{F}^{\beta}}$ instead, see also Remark 1, point 5.
$10^{3}$$10^{4}$$10^{5}$111122223333$|\mathcal{T}_{h}|$#$\#its,\epsilon=10^{-5}$$\gamma=10$$\gamma=10^{3}$$\gamma=10^{6}$$10^{3}$$10^{4}$$10^{5}$111122223333$|\mathcal{T}_{h}|$#$\#its,\epsilon=1$
Figure 1: Number of iterations for the example of Section 6.3 for different stabilization factors $\gamma$ and with $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$.
The choice of $\beta$:
Although the theory requires that $\beta\geq 4$, we discuss in the following the convergence for different (smaller) choices $\beta=2,3,4$. We choose $tol_{n}=10^{-9}$ and set $\gamma=10$. Furthermore, in contrast to the previous example, we choose $\varepsilon=10^{-3}$, as our computations showed faster pre-asymptotic convergence of some errors for smaller choices. Consequently, we increased the diffusion coefficient to make the jumps more pronounced. In Table 3 we again present the error $\|\nabla(u-u_{h}^{+})\|_{L^{2}(\Omega)}$ but further present the values of the jump norm $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}=\|[\![(u_{h}^{+})^{0}]\!]\|_{\mathcal{F}_{h}}$ and the $L^{2}$-norm $\|(u_{h}^{+})^{0}\|_{L^{2}(\Omega)}$. We observe that the error $\|\nabla(u-u_{h}^{+})\|_{L^{2}(\Omega)}$ is independent of the choice of $\beta$ and converges with the optimal order. Note that the same conclusions can be made for the error $\|(u-u_{h}^{+})\|_{L^{2}(\Omega)}$, which is omitted here for brevity. The jump norm $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ shows a faster convergence. This suggests that the convergence analysis presented in Corollary 1 may be sharpened. More precisely, the numerical results indicate that $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ converges at a rate of order $h^{\beta}$. As in the previous paragraph, this observation can be motivated by the fact that for $\beta=1$, the stabilization essentially reduces to a standard interior penalty method, for which linear convergence of order $\mathcal{O}(h)$ is typically expected.
Since standard error estimates are often based on ellipticity with respect to $a_{h}$, as used in the first step of the proof of Lemma 9, we can argue that for $\beta>1$
$\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}^{2}=\|[\![u_{h}^{0}]\!]\|_{\mathcal{F}_{h}}^{2}\cdot\frac{h^{\beta-1}}{h^{\beta-1}}\lesssim h^{\beta-1}a_{h}(u_{h}^{0},u_{h}^{0})\lesssim h^{\beta-1}\cdot\mathcal{O}(h)=\mathcal{O}(h^{\beta}).$
$\beta=1$
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $3.16\cdot 10^{-4}$ (--- -) $3.49\cdot 10^{-4}$ (--- -)
$1984$ $1.51\cdot 10^{-1}$ ($1.04$) $9.84\cdot 10^{-5}$ ($1.68$) $2.01\cdot 10^{-4}$ ($0.80$)
$7936$ $7.49\cdot 10^{-2}$ ($1.01$) $2.85\cdot 10^{-5}$ ($1.79$) $1.25\cdot 10^{-4}$ ($0.69$)
$31744$ $3.73\cdot 10^{-2}$ ($1.01$) $8.19\cdot 10^{-6}$ ($1.80$) $6.93\cdot 10^{-5}$ ($0.85$)
$126976$ $1.86\cdot 10^{-2}$ ($1.00$) $2.10\cdot 10^{-6}$ ($1.96$) $3.58\cdot 10^{-5}$ ($0.95$)
$\beta=2$
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $4.03\cdot 10^{-5}$ (--- -) $3.23\cdot 10^{-5}$ (--- -)
$1984$ $1.51\cdot 10^{-1}$ ($1.04$) $8.64\cdot 10^{-6}$ ($2.22$) $9.10\cdot 10^{-6}$ ($1.83$)
$7936$ $7.49\cdot 10^{-2}$ ($1.01$) $1.36\cdot 10^{-6}$ ($2.67$) $2.79\cdot 10^{-6}$ ($1.70$)
$31744$ $3.73\cdot 10^{-2}$ ($1.01$) $1.83\cdot 10^{-7}$ ($2.89$) $7.77\cdot 10^{-7}$ ($1.85$)
$126976$ $1.86\cdot 10^{-2}$ ($1.00$) $2.03\cdot 10^{-8}$ ($3.17$) $2.01\cdot 10^{-7}$ ($1.95$)
$\beta=3$
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $3.47\cdot 10^{-6}$ (--- -) $2.88\cdot 10^{-6}$ (--- -)
$1984$ $1.51\cdot 10^{-1}$ ($1.04$) $3.82\cdot 10^{-7}$ ($3.19$) $4.10\cdot 10^{-7}$ ($2.81$)
$7936$ $7.49\cdot 10^{-2}$ ($1.01$) $3.01\cdot 10^{-8}$ ($3.67$) $6.35\cdot 10^{-8}$ ($2.69$)
$31744$ $3.73\cdot 10^{-2}$ ($1.01$) $2.01\cdot 10^{-9}$ ($3.91$) $8.87\cdot 10^{-9}$ ($2.84$)
$126976$ $1.86\cdot 10^{-2}$ ($1.00$) $1.11\cdot 10^{-10}$ ($4.18$) $1.15\cdot 10^{-9}$ ($2.95$)
$\beta=4$
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $2.87\cdot 10^{-7}$ (--- -) $2.60\cdot 10^{-7}$ (--- -)
$1984$ $1.51\cdot 10^{-1}$ ($1.04$) $1.59\cdot 10^{-8}$ ($4.17$) $1.88\cdot 10^{-8}$ ($3.79$)
$7936$ $7.49\cdot 10^{-2}$ ($1.01$) $6.36\cdot 10^{-10}$ ($4.64$) $1.47\cdot 10^{-9}$ ($3.68$)
$31744$ $3.73\cdot 10^{-2}$ ($1.01$) $2.15\cdot 10^{-11}$ ($4.88$) $1.03\cdot 10^{-10}$ ($3.83$)
$126976$ $1.86\cdot 10^{-2}$ ($1.00$) $6.06\cdot 10^{-13}$ ($5.15$) $6.70\cdot 10^{-12}$ ($3.95$)
Table 3: Error convergence for the example of Section 6.3 with the diffusion coefficient $\varepsilon=10^{-3}$, stabilization $\gamma=10$ and tolerances $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$ and varying $\beta=1,2,3,4$.
6.4 Interior layer example
We consider the domain $\Omega=(0,1)^{2}$ and choose the parameters $\varepsilon=10^{-7}$ and $\mu=1$. This time, the right-hand side is given by
$f=\begin{cases}0&\text{ in }[\tfrac{1}{4},\tfrac{3}{4}]^{2},\\ 1&\text{ else}.\end{cases}$
Due to the discontinuous right-hand side and the small parameter $\varepsilon$ we expect that the solution will exhibit an interior layer.
To investigate this in more detail, Figure 2 shows the solution of the standard EG method and the solution $u_{h}^{+}$ of our proposed method for different stabilization parameters. Herein, we show results obtained using $\gamma=10$, but larger values of $\gamma$ give very similar results. For the standard method, we directly solved the problem (i.e., without using an iterative scheme) and employed $\beta=1$, $\gamma=10$, and $\alpha=0$ (no stabilization by (10)). The computations were made on a mesh with $|{\mathcal{T}_{h}}|=242$ and the same tolerances and damping parameters as in the previous section. We observe that the solution of the standard EG method exhibits a highly oscillatory behavior in the interior layer and fails to preserve the limits. In contrast, the solution of our proposed method is bound-preserving. We want to emphasize that, although the solution appears to be approximated solely by the linear Lagrange Finite element - similarly to [BGPV24] - our method is locally conservative since the piecewise constant part of $u_{h}^{+}$ is very small thanks to the over-penalization.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.2
Figure 2: Solutions of the interior layer example of Section 6.4 for the standard EG method (left) and the proposed method with $\gamma=10$ (right). For better visualization, the solution was cut along the line $[0,1]\times\{0.5\}$.
7 Conclusion
We have proposed a bound-preserving EG method whose solution is locally (and globally) conservative. A fundamental tool to achieve this is the way we have built the limiting process, by leaving the piecewise constant part free, and limiting the piecewise linear part in such a way that the sum respects the bounds given by the continuous problem. As a result, the stabilization needed to compensate for the ”unconstrained” part of the solution does not depend on the piecewise constant part of it, which implies conservation. In addition, it is important to insist on the fact that, although for the analysis a variational inequality was used, the method itself is not equivalent to a variational inequality (unlike [BGPV24, ABP24, BPT25]). The method is proven to to approximate smooth solutions with optimal convergence rates if the analytical solution respects pre-defined upper and lower bounds. If the solution is not smooth, our approach still respects the upper and lower bounds, is mass conservative, and convergent, but the significance of the penalty parameter $\gamma$ increases: a larger $\gamma$ suppresses oscillations stronger than a smaller $\gamma$. However, the nonlinear problem remains solvable independent of our tuning parameters $\gamma$ and $\beta$ (which we fix to $4$). In addition, the use of the splitting algorithm presented in Section 6.1 allows us to completely bypass the ill-conditioning of the linear systems that would arise were (9) to be solved in a monolithic way.
Thus, we have provided an efficient numerical framework to approximate diffusion–reaction problems. Naturally, some open problems should be tackled in future research. Some of these challenges include:
  • An adaptive strategy to select $\beta$ and $\gamma$ in the spirit of Section 4.4.
  • The consideration of non-linear and hyperbolic equations.
  • The open problem whether the super-convergence of the $L^{2}$ error can be explained by a more involved proof strategy that allows for mimicking the Aubin–Nitsche trick.